Atlantic Tires is considering an investment in a new production facility. For the next five years, the company expects to achieve an annual cash flow of $3,500,000 if the project is successful and $1,000,000 if the project is a failure. Projects of this nature are successful 60% of the time. The initial investment required is $9,500,000, and the relevant discount rate is 10%. Assume the production facility will be obsolete five years from now (i) What is the base-case NPV? (ii) Suppose the company has an option to double the scale of the project in one year if the project is a success at the end of the first year. This would imply that if the project is successful, project annual cash flows would be $7,000,000 after expansion. What is the NPV of the project? Consider the possibility of expansion in answering (iii) what is the value of the option to expand? In the previous example, assume abandonment is not an option if the project is a failure and consider the followin (i) Based on the first years performance, the project can be dismantled and sold for $4,500,000 at the end of the first year. What is the NPV of the project? Consider the possibility of abandonment in answering (i) What is the value of the option to abandon? In the previous example, assume abandonment is not an option if the project is a failure. In addition, suppose the cost of the new production facility will decline by $1,900,000 next year Should the company make the investment today or delay it by one year?
Answer #1

1. Base Case NPV

Initial Investment Required = $9,500,000

Discount Rate = 10%

Project Tenor = 5 years

Annual Cash Flow In case of Success = $3,500,000

Annual Cash Flow in case of Failure = $1,000,000

Success Rate = 60%

Success Rate adjusted yearly cash flow = 0.6 * 3,500,000 + 0.4 * 1,000,000

Success Rate adjusted yearly cash flow = 2,100,000 + 400,000 = 2,500,000

PV =-I t + X Cash flowt/ (1 + r)t natial Investmen

NPV = -$23,033.08

2.

Incase of success and doubling of scale of project after Year 1

Cash Flow of Year 1 = 3,500,000 - 9,500,000 = -$6,000,000

Doubling of scale will lead to doubling of investment at Year 1

Cash Flow from Year 2 to Year 5 = $7,000,000

NPV = $5,217,325.57

3.

Value of option of expansion = NPV in success with expansion - NPV in success without expansion

NPV in success without expansion = $3,767,753.69

Value of option of expansion = 5,217,325.57 - 3,767,753.69

Value of option of expansion = $1,449,571.88